Liquid hexane will react with gaseous oxygen to produce gaseous carbon dioxide and gaseous water . Suppose 83. g of hexane is mixed with 100. g of oxygen. Calculate the minimum mass of hexane that could be left over by the chemical reaction. Be sure your answer has the correct number of significant digits.
Answer #1

Given, Liquid Hexane react with gaseous oxygen to produce goseous co, and 9egeous H, О. Co, (9) + Нехоо (2) + ox930ч () — Но(Molar mass of Hexane = 6x atomic mass of Carbon + lux atanic mass of Hydrogen = 6х12 + (4х1 = 72 +1° = 8% 3]що , Molar mass o